During the first week of January, an employee works 48 hours. For this company, workers earn 150% of their regular rate for hours in excess of 40 per week. Her pay rate is $10 per hour, and her wages are subject to no deductions other than FICA Social Security, FICA Medicare, and federal income taxes. The tax rate for Social Security is 6.2% of the first $128,400 earned each calendar year and the FICA tax rate for Medicare is 1.45% of all earnings. The current FUTA tax rate is 0.6%, and the SUTA tax rate is 5.4%. Both unemployment taxes are applied to the first $7,000 of an employee’s pay. The employee has $87 in federal income taxes withheld. What is the amount of this employee’s net pay for the first week of January? (Round your intermediate calculations to two decimal places.)
Answer #1
Gross earnings 520.00 =(40*10)+(8*10*150%)
Less: Deductions
Social Security tax 32.24 =520*6.2%
Medicare tax 7.54 =520*1.45%
Federal income tax 87.00
Total Deductions 126.78
Net pay for the first week of January 393.22